OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to give line express in y=mx+b (5,-2), 2x+9y=5

OpenStudy (mimi_x3):

m = 2/9 -2 = 2/9*5 +b -2 = 10/9 + b b = -28/9 Therefore , y + 2/9x-28/9 (not sure If im right though , someone should check it xD)

OpenStudy (saifoo.khan):

Let's check.

OpenStudy (mimi_x3):

saifoo am i right ?

OpenStudy (saifoo.khan):

not 2/9

OpenStudy (mimi_x3):

woops my bad didnt read the question properly xD

jimthompson5910 (jim_thompson5910):

slope should be -2/9

OpenStudy (saifoo.khan):

O_O Jim's right.

OpenStudy (mimi_x3):

LOL, saifoo's wrong for the first time xD

OpenStudy (saifoo.khan):

haha, im a Human too! :p

OpenStudy (saifoo.khan):

Solve it now.

OpenStudy (mimi_x3):

LOL y = -2/9 - 8/9 Am i right , right >?

OpenStudy (mimi_x3):

y = -2/9x - 8/9 *

OpenStudy (mimi_x3):

wtf , how is that possible >

OpenStudy (saifoo.khan):

why isn't it?

OpenStudy (mimi_x3):

isnt it -2/9 xD

OpenStudy (mimi_x3):

i didnt understand what you wrote before xD

OpenStudy (saifoo.khan):

\[y = -\frac{2 x}{9}-\frac{8}{9 }\]

OpenStudy (mimi_x3):

LOL , i was right xD

OpenStudy (saifoo.khan):

The fact is, i never said u r wrong.

OpenStudy (mimi_x3):

lols ok

OpenStudy (saifoo.khan):

HAHA. u can look at this thread.

OpenStudy (anonymous):

so what would it be?

OpenStudy (saifoo.khan):

mine.

OpenStudy (saifoo.khan):

the one in bold letters. and fractions.

OpenStudy (mimi_x3):

my one is right as well xD , can you look at this thread ?